# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v2

@@ Abstract
Satellite communications present challenging characteristics. Continuous end-to-end connectivity may not be available due to the large distances between satellites. Moreover, resources such as link capacity and buffer memory may be limited. Routing in satellite networks is therefore both complex and crucial to avoid packet losses and long delays. The Delay Tolerant Network (DTN) paradigm has emerged as an efficient solution for managing these challenging networks. Contact Graph Routing (CGR), a deterministic routing algorithm, is one of the most popular DTN algorithms. CGR is compatible with the “store, carry, and forward” principle, whereby a node receives a message and stores it in its buffer until a transmission opportunity becomes available. However, CGR relies on simplified models to incorporate potential constraints in the route search. For instance, the linear volume assumption is often used to consider capacity constraints. Moreover, capacity management and buffer management are mostly performed during the forwarding phase, once an issue has occurred. In this paper, we propose taking measures before or during the route search in order to find routes that respect both contact-capacity and node-buffer limits. We introduce the contact splitting and edge pruning operations to effectively account for the routing constraints. This ensures that CGR outputs the optimal solution among the subset of valid solutions. The proposed approach can also be used to book resources to be used in case of issues during the forwarding phase.
Index Terms:
Delay tolerant networks, contact graph routing, satellite communications, buffer, contact capacity.
@@ I Introduction
As a first networking solution for satellite systems, traditional Internet Protocol (IP)-based routing algorithms have been adapted for satellite scenarios where the network topology can be assumed stable and where the connectivity can be considered continuous. These IP-based algorithms require explicit topology-update notifications [3][4]. In case of frequent network-topology changes, this approach therefore results in significant overhead which can drain the limited storage and processing resources of satellites. Moreover, these IP-based algorithms are designed for scenarios where end-to-end connectivity is available. However, satellite networks are often characterized by intermittent connectivity. This intermittent nature induces multiple discontinuous episodes of communication opportunities between pairs of nodes (so-called contacts, see the following sections). Limited effectiveness is therefore observed in non-dense satellite scenarios, such as the initial-stage deployment of a constellation.
@@
Moreover, in addition to the delivery-time objective, optimizing the delivery rate and the resource consumption should also be considered. For instance, the use of transmission resources must be carefully optimized because of their limited availability. Additionally, as data units are often held in the node buffers for long periods, optimizing storage resources to avoid congestion and loss is crucial. Nevertheless, the current version of CGR and the related latest research do not fully consider or fail to optimize contact capacity and buffer occupancy. Instead, they only provide local and reactive management during the forwarding phase. This leads to potential resource waste, sub-optimal solutions, and delays, especially in a resource-constrained environment. It is therefore important to explore alternative approaches.
One modern application where capacity and buffer management may be critical is cislunar communications. With the advent of the Artemis Program [2], many projects in the lunar environment have arisen, triggering a high need for communication resources [16, 17]. Another application is low-earth orbit satellite constellations, which are expected to be utilized extensively in the future.
@@
Main contributions. We provide a routing solution based on CGR with a global and proactive congestion control. We propose to ensure contact-capacity management and buffer-limit management before or during the route search such that the obtained route is optimal and respects these two constraints. The objective is achieved as follows.
@@
The modified network information enables the removal of routes that do not respect the constraints (and only these routes). Subsequently, the route-search algorithm, such as CGR, can find the optimal solution among the subset of valid solutions.
@@
Structure of the paper. The paper is structured as follows. In Section II, we first introduce key terms used in the paper and we provide a review of CGR. In Section III, we focus on existing capacity and buffer management challenges in DTNs to introduce our motivations. The proposed solution is described in Section IV. The source-routing CGR benchmark is detailed in Section V. Simulation assumptions and results comparing our method with the benchmark are given in Sections VI and VII, respectively.
@@ II-A Key notions
  • Node: In a satellite communication scenario, a node $X$ can be a satellite or a Ground Station (GS) as illustrated in Figure 1.
@@
  • Contact: A contact $\text{C}_{j}$ is defined by a communication opportunity between two nodes. It is characterized by sender and receiver nodes, a start time $t_{1,j}$, an end time $t_{2,j}$, and a data rate/capacity $R_{\text{C}_{j}}$ in bits per second (bps). The volume of a contact is its rate multiplied by its duration.
@@
  • Contact Plan (CP): A CP is the list of contacts for a time period. It reflects the evolution of the network topology over time. An example is provided in Table I.
  • Edge: An edge exists between two contacts when the receiver of the first contact matches the sender of the second contact, and the end time of the first contact is earlier than the start time of the second.
@@
  • Bundle: A bundle is the data unit encoded by the BP. A bundle $i$ is characterized by its size $S_{i}$ in bits. A bundle is generated at a source node and is assigned a destination node.
@@
  • Route: A route or a path is a series of contacts that are used to deliver a bundle from its source node to its destination node.
@@ II-B Contact graph routing (history)
Since its invention, CGR has gained attention in the research community and several improvements have been proposed. Using Dijkstra’s algorithm for the route search and choosing the Best-case Delivery Time (BDT) as objective were presented in [18]. Then, the Earliest Transmission Opportunity (ETO) was introduced to compute the time at which a bundle is transmitted through the contact [19]. The ETO takes into account the time a bundle has to wait before being transmitted because of other scheduled bundle transmissions on the same contact. Later, a complementary overbooking management was added to the ETO to improve routing decisions [20]. Nevertheless, these managements are performed during the forwarding phase. Furthermore, a route list-management technique using Yen’s algorithm was proposed to further enhance the standard CGR algorithm [21]. Note however that this route management technique increases processing complexity during scheduling [22]. These enhancements among others result in the currently used version of CGR [13].
@@
Research continues to improve the latest version of CGR, focusing on the optimization of limited resources including energy [23], computation [24, 25], buffer memory [26], and link capacity [27].
II-C Contact graph routing (implementation)
@@ II-C1 CP construction
II-C2 Route search
Route search is performed using an adapted version of Dijkstra’s algorithm [29]. This adaptation considers the discontinuities of Intermittently Connected Networks (ICNs) by using a contact-graph representation of the network rather than a node-graph representation. The output of a Dijkstra search is the route, consisting of a succession of contacts from the CP, connecting the source to the destination with the lowest BDT.
@@ II-C2 Route search
Algorithm 1 presents both the original CRP and our contributions. Our modifications are marked with asterisks (**, lines $2$ and $8-12$) in the algorithm. In this section, we explain only the original CRP. The modifications will be explained in Section IV-B2.
@@
  • A successor of a contact. I.e., the conctact that comes after in a route.
  • A predecessor of a contact. I.e., the contact that comes before in a route.
  • An arrival time $t_{\text{arr},i}$ that indicates the time needed to reach the receiver node of the contact.
@@
A contact $\text{C}_{i}\in$ CP must meet the following criteria to be explored (otherwise it is ignored) (lines $5-7$). If these criteria are met $C_{i}$ will be considered as a potential successor of $C_{s}$.
a) $\text{C}_{i}$ must have the selected contact’s receiver as its sender.
b) Let $t_{\text{arr},i}$ be the arrival time at $\text{C}_{i}$, i.e., the time at which the receiver node of contact $\text{C}_{i}$ is reached through the succession of contacts that led to $\text{C}_{i}$. The end time of $\text{C}_{i}$ must be greater than the arrival time at the receiver of the selected contact, $t_{\text{arr},s}$. This means that the successor $\text{C}_{i}$ must not end before the bundle arrives at its sender node (the receiver node of $\text{C}_{s}$).
@@
f) The remaining volume of $\text{C}_{i}$ must be greater than zero. This means that after allocating $\text{C}_{i}$ to bundles and deducting bundle size from the volume of $\text{C}_{i}$, some volume must remain available in $\text{C}_{i}$ to allow more data to pass through.
Once a contact has passed all checks, it is considered as an explored contact. The arrival time of an explored contact is computed according to the time taken to reach the receiver of the explored contact following the route that led to it (line $13$). If the computed arrival time is less than its last calculated value $t_{\text{arr},i}$ (in the case where $\text{C}_{i}$ has already been explored), it is updated (lines $14$ and $15$). In addition, the selected contact $\text{C}_{s}$ becomes the predecessor of $\text{C}_{i}$ and the visited nodes of $\text{C}_{i}$ are updated by adding the visited nodes of the nodes within the succession of contacts that led to $\text{C}_{i}$ (lines $16$ and $17$) (to avoid getting back to them, which would induce loops).
Each time the destination node is reached (i.e., when the receiver of an explored $C_{i}$ is the bundle’s destination), the arrival time is saved as BDT if it is lower than the previously calculated BDT. In such cases, $\text{C}_{i}$ is considered as a final contact $\text{C}_{\text{end}}$ with its corresponding BDT (lines $18-21$). After finishing all contacts in the CP, the currently selected contact $\text{C}_{s}$ is marked as visited to avoid repeating the same work later (line $24$).
@@
2:** Additional input: Bundle b, $T_{1,s}=\{t_{1,o_{l}}\}\mid o_{l}\in\mathcal{O}_{Z_{1}}$, $T_{2,s}=\{t_{2,o_{l}}\}\mid o_{l}\in\mathcal{O}_{Z_{1}}$, where ${Z_{1}}$ is the receiver of $\text{C}_{s}$*
@@
5:     if sender of $\text{C}_{i}\neq$ receiver of $\text{C}_{s}$ or $t_{i,2}\leq t_{\text{arr},s}$ or $\text{C}_{i}$ is already visited or receiver of $\text{C}_{i}$ is a visited node by $\text{C}_{s}$ or $\text{C}_{i}$ is suppressed or remaining volume of $\text{C}_{i}==0$ then
@@
9:     $T_{1,s}=[T_{1,s},\infty]$
10:     $T_{2,s}=[0,T_{2,s}]$
11:     if any ($t_{\text{arr},s}\geq T_{2,s}$ and $t_{i,1}\geq T_{1,s}$then
@@
15:     compute $t_{\text{arr}}$ to the receiver of $\text{C}_{i}$, which follows $\text{C}_{s}$
16:     if $t_{\text{arr}}<t_{\text{arr},i}$ then
17:         $t_{\text{arr},i}=t_{\text{arr}}$
@@
19:         $\text{C}_{i}$’s visited nodes $=\text{C}_{s}$’s visited nodes and $\text{C}_{i}$’s     receiver
20:         if $\text{C}_{i}$’s receiver is bundle’s destination and     $t_{\text{arr},i}<$ BDT then
21:              BDT $=t_{\text{arr},i}$

### v3

@@ Abstract
Satellite communications present challenging characteristics. Continuous end-to-end connectivity may not be available due to the large distances between satellites. Moreover, resources such as link capacity and buffer memory may be limited. Routing in satellite networks is therefore both complex and crucial to avoid packet losses and long delays. The Delay Tolerant Network (DTN) paradigm has emerged as an efficient solution for managing these challenging networks. Contact Graph Routing (CGR), a deterministic routing algorithm, is one of the most popular DTN algorithms. CGR is compatible with the “store, carry, and forward” principle, whereby a node receives a message and stores it in its buffer until a transmission opportunity becomes available. However, CGR relies on simplified models to incorporate potential constraints in the route search. For instance, the linear volume assumption is often used to consider capacity constraints. Moreover, capacity management and buffer management are mostly performed during the forwarding phase, once an issue has occurred. These reactive management techniques cause many collisions and increase the average delivery time. In this paper, we propose taking measures before or during the route search in order to find routes that respect both contact-capacity and node-buffer limits. We introduce the contact splitting and edge pruning operations to effectively account for the routing constraints. This ensures that CGR outputs the optimal solution in terms of delivery time among the subset of valid solutions. The problem is formalized as the Feasible Earliest-Arrival Path with Capacity and Buffer constraints (FEAP-CB) and optimality is proved. The proposed approach can also be used to book resources to be used in case of issues during the forwarding phase.
@@ I Introduction
As a first networking solution for satellite systems, traditional Internet Protocol (IP)-based routing algorithms have been adapted for satellite scenarios where the network topology can be assumed stable and where the connectivity can be considered continuous. These IP-based algorithms require explicit topology-update notifications [1][31]. In case of frequent network-topology changes, this approach therefore results in significant overhead which can drain the limited storage and processing resources of satellites. Moreover, these IP-based algorithms are designed for scenarios where end-to-end connectivity is available. However, satellite networks are often characterized by intermittent connectivity. This intermittent nature induces multiple discontinuous episodes of communication opportunities between pairs of nodes (so-called contacts, see Section II-A). Limited effectiveness is therefore observed in non-dense satellite scenarios, such as the initial-stage deployment of a constellation.
@@
Moreover, in addition to the delivery-time objective, optimizing the delivery rate and the resource consumption should also be considered. For instance, the use of transmission resources must be carefully optimized because of their limited availability. Additionally, as data units are often held in the node buffers for long periods, optimizing storage resources to avoid congestion and loss is crucial. Nevertheless, the current version of CGR and the related latest research do not fully consider or fail to optimize contact capacity and buffer occupancy. Instead, they only provide local and reactive management during the forwarding phase. This leads to potential resource waste, delays, errors, and more generally sub-optimal solutions which may be harmful especially in a resource-constrained environment (see also Section III-A for more details on this aspect). It is therefore important to explore alternative approaches.
One modern application where capacity and buffer management may be critical is cislunar communications. With the advent of the Artemis Program [3], many projects in the lunar environment have arisen, triggering a high need for communication resources [42, 61]. Another application is low-earth orbit satellite constellations, which are expected to be utilized extensively in the future. Moreover, the IOAG Mars &\& Beyond Communications (MBC) architecture explicitly plans the infusion of DTN-based network management to orchestrate relay orbiters, surface hubs, and Earth stations.[54][27].
@@
Main contributions. We provide a routing solution based on CGR with a global and proactive congestion control. We propose to ensure contact-capacity management and buffer-limit management before or during the route search such that the obtained route is optimal and respects these two constraints. Optimality means achieving the lowest possible delivery time. The formal optimization problem (FEAP-CB) is given in Section III, and the optimality proof in Section IV-C. The objective is achieved as follows.
@@
The modified network information enables the removal of routes that do not respect the constraints (and only these routes). Subsequently, the route-search algorithm, such as CGR, can find the optimal solution among the subset of valid solutions. It means that the algorithm will automatically find the optimal storage location of bundles (from delivery time perspective) considering the forecasted buffer status.
@@
Assumptions in this paper. This study proposes a solution in an ideal scenario. We have the main three assumptions:
  • We assume instantaneous sharing of the network information among source routing nodes, perfect topology knowledge, no unexpected disruptions.
  • We do sequential per packet optimization not full traffic optimization.
  • Complexity is not the main concern.
Under these assumptions, we show that the proposed solution avoids all routing errors, both theoretically and with simulations. We acknowledge that these assumptions are not realistic. This work can therefore be understood as a theoretical foundation under ideal assumptions and may serve as a benchmark in the future for suboptimal solutions operating with partial knowledge, such as Machine Learning (ML)-based approaches.
Nevertheless, note that [15, Sec. IV] also investigates the notion of contact segments (conceptually close to contact splitting, proposed at the same period) but considers the complementary simulations in a distributed setting with local, unsynchronized volume views (i.e., without assuming a globally shared, up‑to‑date CP volume state). Simulations show the advantage also in this case, see Section II-C1 for more details.
Structure of the paper. The paper is structured as follows. In Section II, we first introduce key terms used in the paper and we provide a review of CGR. In Section III, we focus on existing capacity and buffer management challenges in DTNs to introduce our motivations and we present the mathematical problem formulation (FEAP-CB). The proposed solution is described in Section IV, which also includes a proof of optimality of the approach. The source-routing CGR benchmark is detailed in Section V. Simulation assumptions and results comparing our method with the benchmark are given in Sections VI and VII, respectively. Symbols used in the paper are summarized in the Appendix (Table IV).
@@ II-A Key notions
  • Node: In a satellite communication scenario, a node can be a satellite or a Ground Station (GS) as illustrated in Figure 1. The set of all nodes is $\mathcal{N}$. In the text we denote a node by $n$ or by context-specific letters ($s$, $d$ for source and destination of a bundle, $u_{i}$, $v_{i}$ for sender and receiver of contact $\text{C}_{i}$). In the figures we use capital letters ($X$, AA, $B$, etc.) for clarity.
@@
  • Contact: A contact $\text{C}_{j}$ is defined by a communication opportunity between two nodes. It is characterized by sender and receiver nodes, a start time $t_{1,j}$, an end time $t_{2,j}$, and a data rate/capacity $R_{j}$ in bits per second (bps). The volume of a contact is its rate multiplied by its duration. Each contact $\text{C}_{i}$ is therefore a tuple
    $\begin{split}&\text{C}_{i}=(u_{i},\,v_{i},\,t_{1,i},\,t_{2,i},\,R_{i},\,\delta_{i}),\quad\\ &t_{1,i}<t_{2,i},\ \ R_{i}>0,\ \ \delta_{i}\geq 0,\end{split}$ (1)
    where $\delta_{i}\geq 0$ is the one-way propagation delay.
@@
  • Contact Plan (CP): A CP is the list of contacts for a time period. It reflects the evolution of the network topology over time. An example is provided in Table I. The set of contacts is either denoted by $V_{c}^{0}$ or simply CP.
@@
  • Bundle: A bundle is the data unit encoded by the BP. A bundle $i$ is characterized by its size $S_{i}$ in bits. A bundle of size $S>0$ is released at source $s\in\mathcal{N}$ at time $t_{0}$ and must reach destination $d\in\mathcal{N}$.
  • Edge: An edge $(i\to j)$ exists between two contacts when the receiver of the first contact $i$ matches the sender of the second contact $j$ ($v_{i}=u_{j}$). $E_{c}^{0}=\{(i\to j):v_{i}=u_{j}\}$ is the set of such edges, for all contacts in $V_{c}^{0}$.
    For a given bundle of size $S$, the feasible window of contact $\text{C}_{i}$ is $W_{i}=[t_{1,i},\,t_{2,i}-S/R_{i}]$: the set of start times at which a transmission of $S$ bits on $\text{C}_{i}$ finishes by $t_{2,i}$. Thus each transmission must finish before the contact ends, and each next hop must start no earlier than the arrival $f_{i}=\tau_{i}+S/R_{i}+\delta_{i}$ from the previous hop. For a given bundle of size $S$, only a subset of $E_{c}^{0}$ that satisfy these timing constraints is actually usable [^1] An equivalent modeling is to enforce this condition and the schedulability directly in the contact graph by reducing $E_{c}^{0}$..
  • Buffer capacity. For each node $n$, let $B_{n}(t)$ denote the (full) buffer capacity (in bits) available at time $t$. In this paper, the buffer capacity is constant over time, so $B_{n}(t)=B_{\text{max}}$ for all $t$. Let $Q_{n}(t)$ be the booked buffer load due to previously scheduled bundles ($Q_{n}(t)=0$ if no prior bookings). We define the residual buffer capacity as
    $\bar{B}_{n}(t)=\max\{\,0,\ B_{n}(t)-Q_{n}(t)\,\}.$
  • Connectivity-only contact graph $(V_{c}^{0},E_{c}^{0})$: We denote by $(V_{c}^{0},E_{c}^{0})$ the connectivity-only contact graph before any splitting: vertices are the contacts in the CP, and edges are as in the ones described in the paragraph above.
  • Feasible contact graph $(V_{c},E_{c})$: Contact splitting (see Section IV-A1) turns $V_{c}^{0}$ into the vertex set $V_{c}$ of split contacts. On $V_{c}$, the connectivity-only edge set is again denoted $E_{c}^{0}$. The feasible contact graph $(V_{c},E_{c})$ is obtained by removing additional edges from $E_{c}^{0}$ via edge pruning (see Section IV-A2), so $E_{c}\subseteq E_{c}^{0}$. How $(V_{c},E_{c})$ is built is detailed in Section IV.
@@
  • Route (path) and schedulability: A route (or path) is a sequence of contacts $P=(\text{C}_{1},\ldots,\text{C}_{k})$ used to deliver a bundle from source $s$ to destination $d$, with the basic connectivity constraint that the receiver of $\text{C}_{i}$ equals the sender of $\text{C}_{i+1}$:
    $u_{1}=s,\quad v_{k}=d,\quad v_{i}=u_{i+1}\ \ (i=1,\ldots,k-1).$
    In the connectivity-only contact graph $(V_{c}^{0},E_{c}^{0})$, this means $\text{C}_{i}\in V_{c}^{0}$ and $(\text{C}_{i},\text{C}_{i+1})\in E_{c}^{0}$ for all $i$, so $P$ is a path in $(V_{c}^{0},E_{c}^{0})$.
    When considering transmission timing for a given bundle, a path $P=(\text{C}_{1},\ldots,\text{C}_{k})$ is schedulable if it satisfies the timing constraints (feasible windows and precedence) introduced in the Edge paragraph above. This schedulability is enforced in the path selected by CGR when the route search runs on the contact graph: the (time-dependent) Dijkstra search explores only contact successions that respect these constraints.
  • Feasible route: A feasible route is a path (in the connectivity graph $(V_{c}^{0},E_{c}^{0})$) that is schedulable, contact-capacity feasible, and buffer feasible (see Section III-C). The set of feasible routes for a bundle of size $S$ and release time $t_{0}$ is denoted $\mathcal{P}_{f}(S,t_{0})$. We will show (Lemma 3, Section IV-C) that $\mathcal{P}_{f}(S,t_{0})$ coincides with the set of schedulable paths in the feasible graph $(V_{c},E_{c})$.
@@ II-B Contact graph routing (history)
Since its invention, CGR has gained attention in the research community and several improvements have been proposed. Using Dijkstra’s algorithm for the route search and choosing the Best-case Delivery Time (BDT) as objective were presented in [50]. Then, the Earliest Transmission Opportunity (ETO) was introduced to compute the time at which a bundle is transmitted through the contact [5]. The ETO takes into account the time a bundle has to wait before being transmitted because of other scheduled bundle transmissions on the same contact. Later, a complementary overbooking management was added to the ETO to improve routing decisions [4]. Nevertheless, these management techniques are performed during the forwarding phase. Furthermore, a route list-management technique using Yen’s algorithm was proposed to further enhance the standard CGR algorithm [22]. Note however that this route management technique increases processing complexity during scheduling [14]. These enhancements among others result in the currently used version of CGR [20].
II-C Recent trends in DTN routing and further elements on the SoA
@@
Research continues to improve the latest version of CGR, focusing on the optimization of limited resources including energy [44], computation [21, 43], heterogeneous traffic and buffer-related aspects [40], and link capacity [35].
II-C1 Contact segmentation in [15]
Contact segmentation was proposed independently around the same period as the first version of this paper [15, Sec. IV]. It follows the same idea as contact splitting: instead of splitting the contacts, that approach adds a new parameter to the contacts which describes the available segments of the contact. That work however considers simulations for the distributed case with local, unsynchronized volume views (i.e., without assuming a globally shared, up‑to‑date CP volume state). In [15], the capability of the node is implemented using a new library [16].
The following improvement compared to the benchmark is summarized in the conclusion of [15]: contact segementation delivers up to a 10 % improvement in delivery ratios, reductions in delivery delay of up to 11 % and up to a 16 % decrease in hop counts.
II-C2 Machine-learning routing.
Recent works apply Graph Neural Networks (GNNs) and Deep Reinforcement Learning (DRL) to LEO/DTN routing to adapt under dynamic environments, potentially with partial network knowledge and unexpected events. This is promising for traffic-level optimization where load, not just individual bundles, must be steered and complexity is an issue. However, these methods generally lack shortest-path optimality guarantees and usually handle contact-capacity and buffer limits via rewards or post-hoc checks rather than hard feasibility at selection time [30, 63, 39]. Therefore, current research investigates the trade-off between the theoretical guarantees given by CGR under simplified assumptions and the flexibility brought by ML.
We review some existing works. GAUSS [46] learns to accelerate route selection in scheduled DTNs, reporting up to $2\times$ compute reduction and $3\times$ lower variability versus CGR-style search. However, Section 4.1.2 of that work explicitly notes that this better computational profile comes at the expense of decreased efficacy, with degraded delivery time compared to standard CGR.
Under congestion or schedule errors, learned policies can rebalance load and improve throughput/delay, making ML a practical fallback or online refinement atop deterministic baselines [24, 51]. Other works learn next-hop or path-selection policies and show gains in simulations, e.g., multi-agent DRL adapts under load. DRL +GNN combinations improve aggregate metrics in mega-constellations [38, 24, 51]. GraphSAGE+DQN and graph attention handle topology changes [51, 57].
II-C3 Buffer-aware routing and resource reservation in DTNs.
Many opportunistic DTN schemes have long intertwined routing with buffer management. As mentioned in the introduction, opportunistic refers to networks with stochastic contacts and typically rely on replication/priority and local queue control to mitigate uncertainty [29, 52, 37]. This line of work also considers the dropping policy at the buffer node, which is beyond the scope of this paper.
Within this opportunistic family, MaxProp orders both transmissions and drops (explicit buffer eviction policy) using path-likelihood estimates and global ACKs [9]. RAPID treats routing as a resource allocation problem, replicating bundles in decreasing marginal utility to optimize metrics such as mean delay or deadline-meet ratio [2]. PRoPHET prunes epidemic flooding via delivery predictabilities. Variants such as PRoPHETv2 add energy/buffer awareness in forwarding and queue control [23]. Constrained Spray-and-Wait methods cap replication and incorporate congestion control or dynamic copy allocation to reduce buffer pressure [45, 62, 34]. Beyond replication caps, ORWAR is explicitly resource-aware: it estimates contact windows to avoid partial transmissions and orders replication/drops by message utility (buffer-, time-, and contact-aware scheduling) [48]. This is closer to the approach proposed in this paper, as our routing algorithm considers forecasted buffer resources when choosing the route. Recent surveys emphasize that these approaches largely enforce soft feasibility (via prioritization, dropping, replication limits), whereas our CGR formulation guarantees feasibility during route computation in scheduled DTNs [29, 52, 49].
II-D Contact graph routing (implementation)
@@ II-D1 CP construction
II-D2 Route search
Route search is performed using an adapted version of Dijkstra’s algorithm [19]. This adaptation considers the discontinuities of Intermittently Connected Networks (ICNs) by using a contact-graph representation of the network rather than a node-graph representation. In this standard implementation, the route search runs on the connectivity-only contact graph $(V_{c}^{0},E_{c}^{0})$. In the proposed algorithm (Section IV), it runs on the feasible graph $(V_{c},E_{c})$. The output of a Dijkstra search is the route, consisting of a succession of contacts from the CP, connecting the source to the destination with the lowest BDT.
@@ II-D2 Route search
Algorithm 1 presents both the original CRP and our contributions. Our modifications are marked with asterisks (**, lines $2$ and 8128-12) in the algorithm. In this section, we explain only the original CRP. The modifications will be explained in Section IV-B2.
@@
  • A successor of a contact (the contact that comes after in a route). A successor of $\text{C}_{s}$ is a contact $\text{C}_{i}$ with $(\text{C}_{s},\text{C}_{i})\in E_{c}^{0}$.
  • A predecessor of a contact (the contact that comes before in a route).
  • An arrival time $f_{i}$ at the receiver of the contact. $f_{i}=\tau_{i}+S/R_{i}+\delta_{i}$.
@@
A contact $\text{C}_{i}\in$ CP must meet the following criteria to be explored (otherwise it is ignored) (lines 575-7). If these criteria are met $\text{C}_{i}$ will be considered as a potential successor of $\text{C}_{s}$.
a) $\text{C}_{i}$ must have the selected contact’s receiver as its sender. (Connectivity: $v_{s}=u_{i}$, so $(\text{C}_{s},\text{C}_{i})\in E_{c}^{0}$.)
b) For a bundle of size $S$, $\text{C}_{i}$ is schedulable if $f_{s}\leq t_{2,i}-S/R_{i}$ and $t_{1,i}\leq t_{2,i}-S/R_{i}$, that is, the earliest feasible start $\tau_{i}^{\star}=\max\{t_{1,i},f_{s}\}$ lies in the feasible window $W_{i}=[t_{1,i},\,t_{2,i}-S/R_{i}]$. For readability, Algorithm 1 reference the weaker condition $t_{2,i}\leq f_{s}$.
@@
f) The remaining volume of $\text{C}_{i}$ must be greater than zero. (So the contact has capacity for the bundle, and the feasible window $W_{i}$ is non-empty.) This means that after allocating $\text{C}_{i}$ to bundles and deducting bundle size from the volume of $\text{C}_{i}$, some volume must remain available in $\text{C}_{i}$ to allow more data to pass through.
Once a contact has passed all checks, it is considered as an explored contact. The arrival time of an explored contact is computed according to the time taken to reach the receiver of the explored contact following the route that led to it (line $13$). If the computed arrival time is less than its last calculated value $f_{i}$ (in the case where $\text{C}_{i}$ has already been explored), it is updated (lines $14$ and $15$). In addition, the selected contact $\text{C}_{s}$ becomes the predecessor of $\text{C}_{i}$ and the visited nodes of $\text{C}_{i}$ are updated by adding the visited nodes of the nodes within the succession of contacts that led to $\text{C}_{i}$ (lines $16$ and $17$) (to avoid getting back to them, which would induce loops).
Each time the destination node is reached (i.e., when the receiver of an explored $\text{C}_{i}$ is the bundle’s destination), the arrival time $f_{i}$ is saved as BDT if it is lower than the previously calculated BDT. In such cases, $\text{C}_{i}$ is considered as a final contact $\text{C}_{\text{end}}$ with its corresponding BDT (lines 182118-21). After finishing all contacts in the CP, the currently selected contact $\text{C}_{s}$ is marked as visited to avoid repeating the same work later (line $24$).
@@
2:** Additional input: Bundle b, $T_{1,s}=\{t_{1,o_{l}}\}\mid o_{l}\in\mathcal{O}_{Z_{1}}$ , where $Z_{1}$} is the receiver of $\text{C}_{s}$ **
@@
5:  if sender of $\text{C}_{i}\neq$ receiver of $\text{C}_{s}$ or $t_{2,i}\leq f_{s}$ or $\text{C}_{i}$ is already visited or receiver of $\text{C}_{i}$ is a visited node by $\text{C}_{s}$ or $\text{C}_{i}$ is suppressed or remaining volume of $\text{C}_{i}==0$ then
@@
9:  $\triangleright$ Comment: Compute the first overflow start strictly after $f_{s}$: successors must start before it to avoid buffer overflows.
10:  $a_{\text{next}}\leftarrow\min\{\,t\in T_{1,s}\mid t>f_{s}\,\}$; if no such $t$ then $a_{\text{next}}\leftarrow+\infty$
11:  if $t_{1,i}\geq a_{\text{next}}$ then
@@
15:  compute arrival time $f$ at the receiver of $\text{C}_{i}$ (when extending from $\text{C}_{s}$)
16:  if $f<f_{i}$ then
17:   $f_{i}=f$
@@
19:   $\text{C}_{i}$’s visited nodes $\text{C}_{s}$’s visited nodes and $\text{C}_{i}$’s    receiver
20:   if $\text{C}_{i}$’s receiver is bundle’s destination and    $f_{i}<$ BDT then
21:     BDT $f_{i}$